# GPI-Net: Gestalt-Guided Parallel Interaction Network via Orthogonal Geometric Consistency for Robust Point Cloud Registration

### v1

@@ Abstract
The accurate identification of high-quality correspondences is a prerequisite task in feature-based point cloud registration. However, it is extremely challenging to handle the fusion of local and global features due to feature redundancy and complex spatial relationships. Given that Gestalt principles provide key advantages in analyzing local and global relationships, we propose a novel Gestalt-guided Parallel Interaction Network via orthogonal geometric consistency (GPI-Net) in this paper. It utilizes Gestalt principles to facilitate complementary communication between local and global information. Specifically, we introduce an orthogonal integration strategy to optimally reduce redundant information and generate a more compact global structure for high-quality correspondences. To capture geometric features in correspondences, we leverage a Gestalt Feature Attention (GFA) block through a hybrid utilization of self-attention and cross-attention mechanisms. Furthermore, to facilitate the integration of local detail information into the global structure, we design an innovative Dual-path Multi-Granularity parallel interaction aggregation (DMG) block to promote information exchange across different granularities. Extensive experiments on various challenging tasks demonstrate the superior performance of our proposed GPI-Net in comparison to existing methods. The code will be released at https://github.com/gwk/GPI-Net.
@@ 3.1 Problem Formulation
[…]
@@ 3.5 Dual-Path Multi-Granularity Parallel Interaction Aggregation (DMG)
[…]
@@
[…]
@@ 4 Experiments
RR($\uparrow$) ${}^{\circ}\downarrow$) TE($\downarrow$) IP(%↑) IR(%↑) F1(%↑) Time(s) RR($\uparrow$) ${}^{\circ}\downarrow$) TE($\downarrow$) IP(%↑) IR(%↑) F1(%↑) Time(s)
Traditional SM 55.88 2.94 8.15 47.96 70.69 50.70 0.03 86.57 2.29 7.07 81.44 38.36 48.21 0.03
@@ 4.1 Datasets and Experimental Setup
We assess the effectiveness of GPI-Net through two tasks: outlier rejection and pose estimation. Regarding outlier rejection, we apply inlier precision (IP%), inlier recall (IR%), and F1 score (F1%) as evaluation metrics to measure the accuracy of inliers identification while minimizing false positives. For pose estimation, we adopt registration recall (RR%), rotation error (RE), and translation error (TE). Registration is considered successful when both RE and TE are below the specified thresholds: RE $<$ 15° and TE $<$ 30cm for indoor scenarios, and RE $<$ 5° and TE $<$ 60cm for outdoor scenarios. RE and TE are reported only on successfully registered point clouds.

### v2

@@ Abstract
The accurate identification of high-quality correspondences is a prerequisite task in feature-based point cloud registration. However, it is extremely challenging to handle the fusion of local and global features due to feature redundancy and complex spatial relationships. Given that Gestalt principles provide key advantages in analyzing local and global relationships, we propose a novel Gestalt-guided Parallel Interaction Network via orthogonal geometric consistency (GPI-Net) in this paper. It utilizes Gestalt principles to facilitate complementary communication between local and global information. Specifically, we introduce an orthogonal integration strategy to optimally reduce redundant information and generate a more compact global structure for high-quality correspondences. To capture geometric features in correspondences, we leverage a Gestalt Feature Attention (GFA) block through a hybrid utilization of self-attention and cross-attention mechanisms. Furthermore, to facilitate the integration of local detail information into the global structure, we design an innovative Dual-path Multi-Granularity parallel interaction aggregation (DMG) block to promote information exchange across different granularities. Extensive experiments on various challenging tasks demonstrate the superior performance of our proposed GPI-Net in comparison to existing methods. The code will be released at https://github.com/gwk429/GPI-Net.
@@ 3.1 Problem Formulation
where $\cdot$] is the Iverson bracket, $\delta$ represents a predefined threshold for the inlier, and $\left\|\cdot\right\|$ denotes the Euclidean norm.
@@ 3.5 Dual-Path Multi-Granularity Parallel Interaction Aggregation (DMG)
where $t\in\{1,\cdots,T\}$ and $\textit{Mix\_Bottom\_up}(\cdot)$ consists of an InstanceNorm layer, a BatchNorm layer, a ReLU activation function, and a pointwise convolution layer. The pointwise convolution layer takes input and output features of size $\mathbb{R}^{N\times\mathrm{d}/2^{t-1}}$ and $\mathbb{R}^{N\times\mathrm{d}/2^{t}}$, respectively.
@@
where $t\in\{1,\cdots,T\}$ and $\textit{Mix\_Top\_down}(\cdot)$ consists of an InstanceNorm layer, a BatchNorm layer, a ReLU activation function, and a pointwise convolution layer. The pointwise convolution layer takes input and output features of size $\mathbb{R}^{N\times\mathrm{d}/2^{t+1}}$ and $\mathbb{R}^{N\times\mathrm{d}/2^{t}}$, respectively.
@@ 4 Experiments
[…]
Traditional
SM 55.88 2.94 8.15 47.96 70.69 50.70 0.03 86.57 2.29 7.07 81.44 38.36 48.21 0.03
@@ 4.1 Datasets and Experimental Setup
We assess the effectiveness of GPI-Net through two tasks: outlier rejection and pose estimation. Regarding outlier rejection, we apply inlier precision (IP%), inlier recall (IR%), and F1 score (F1%) as evaluation metrics to measure the accuracy of inliers identification while minimizing false positives. For pose estimation, we adopt registration recall (RR%), rotation error (RE), and translation error (TE). Registration is considered successful when both RE and TE are below the specified thresholds: RE << 15° and TE << 30cm for indoor scenarios, and RE << 5° and TE << 60cm for outdoor scenarios. RE and TE are reported only on successfully registered point clouds.